OpenStudy (anonymous):

Simplify the expression 1/x - 2/y

OpenStudy (anonymous):

Do you have to sole for y?

OpenStudy (anonymous):

*solve

OpenStudy (anonymous):

a/b=c/d ad=bc

OpenStudy (anonymous):

a=1 b=x c=2 d=y Now just substitute.

OpenStudy (anonymous):

It just says to simplify the expression. \[\frac{ 1 }{ x } -\frac{ 2 }{ y }\]

OpenStudy (anonymous):

So 1y = 2x?

OpenStudy (anonymous):

Oh, i was minus! Just make the denominators the same. I have to go, sorry.

OpenStudy (anonymous):

Multiply both by xy